# A Code Knowledge Graph-Enhanced System for LLM-Based Fuzz Driver Generation

### v1

@@ Abstract
The rapid development of large language models (LLMs) with advanced programming capabilities has paved the way for innovative approaches in software testing. Fuzz testing, a cornerstone for improving software reliability and detecting vulnerabilities, often relies on manually written fuzz drivers, limiting scalability and efficiency. To address this challenge, we propose CodeGraphGPT, a novel system that integrates code knowledge graphs with an LLM-powered intelligent agent to automate the fuzz driver generation process. By framing fuzz driver creation as a code generation task, CodeGraphGPT leverages program analysis to construct a knowledge graph of code repositories, where nodes represent code entities, such as functions or files, and edges capture their relationships. This enables the system to generate tailored fuzz drivers and input seeds, resolve compilation errors, and analyze crash reports, all while adapting to specific API usage scenarios. Additionally, querying the knowledge graph helps identify precise testing targets and contextualize the purpose of each fuzz driver within the fuzzing loop. We evaluated CodeGraphGPT on eight open-source software projects, achieving an average improvement of 8.73% in code coverage compared to state-of-the-art methods. Moreover, it reduced the manual workload in crash case analysis by 84.4% and identified 11 real-world bugs, including nine previously unreported ones. This work highlights how integrating LLMs with code knowledge graphs enhances fuzz driver generation, offering an efficient solution for vulnerability detection and software quality improvement.
@@ I Introduction
To address these challenges, we approach fuzz driver generation as a code generation task. By constructing a knowledge graph based on code repositories, we use LLMs to automatically generate fuzz drivers and perform data flow analysis to produce input seeds, thereby achieving fully automated fuzz testing. To enable LLMs to better understand the software under test, we propose abstracting the codebase into a code knowledge graph. In this graph, nodes represent code elements (such as functions, classes, and modules), while edges depict their interactions, with each node containing detailed information about the element’s functionality and behavior. We treat fuzz driver generation as a code planning task, beginning with the formulation of a generation plan followed by iterative fuzz driver generation. To this end, we developed an LLM-driven fuzz driver agent system, CodeGraphGPT, based on the code knowledge graph. Our approach fully automates the fuzzing process, significantly reducing engineering complexity while improving the usability of fuzz testing. The code knowledge graph provides deeper insights into code behavior, guiding the LLM in generating higher-quality API combinations for fuzz drivers and coverage-guided mutation.
We evaluated our approach using eight open-source libraries and addressed three research questions (RQs). First, we compared CodeGraphGPT with other generation techniques (RQ1). Second, we analyzed the influence of each agent in CodeGraphGPT (RQ2). Third, we investigated whether CodeGraphGPT can effectively analyze the causes of crashes generated during fuzzing (RQ3). Our experiments demonstrate that CodeGraphGPT, supported by the code knowledge graph, can generate high-quality fuzz drivers with enhanced code coverage. Through extensive evaluations across these libraries, we found that the agents of CodeGraphGPT play a crucial role in fuzz driver generation. Additionally, our analysis of crash reports indicates that CodeGraphGPT can provide valuable insights for developers.
@@
    We propose CodeGraphGPT, which leverages a code knowledge graph to enhance fuzz driver generation. By analyzing the relationships between code elements and APIs, CodeGraphGPT automatically generates high-quality fuzz drivers that achieve better code coverage.
@@
    CodeGraphGPT introduces a fully automated fuzzing framework that integrates multi-agent systems and LLMs. It automates key steps such as fuzz driver generation, input seed initiation, dynamic program repair, and coverage-guided mutation, significantly reducing manual effort and enhancing overall efficiency in the fuzzing process.
@@
    Experimental results show that CodeGraphGPT achieved an average increase of 8.73% in code coverage compared to the state-of-the-art methods. Furthermore, CodeGraphGPT reduced the manual review workload in crash case analysis by 84.4% and successfully identified 11 real bugs (including nine previously unreported bugs) of tested libraries.
@@ III-A Overview
Figure 1: Overview of CodeGraphGPT
In this section, we present the system design of CodeGraphGPT, a fuzz driver generation framework that integrates a multi-agent system with a code knowledge graph. CodeGraphGPT is designed to enhance the quality and coverage of fuzz testing by efficiently generating fuzz drivers for API combinations. The primary objective of our framework is to harness the rich and relative information within the code knowledge graph to capture the complex invocation relationships among library APIs and, based on this, generate highly effective fuzz drivers. By introducing a coverage-guided mutation strategy, CodeGraphGPT continuously optimizes API combinations, enabling the exploration of additional code paths and the identification of hidden vulnerabilities. Figure 1 illustrates the workflow of CodeGraphGPT, which comprises three main components and a unique feedback loop mechanism.
Initially, CodeGraphGPT parses the target project and its library APIs, extracting and embedding a code knowledge graph from the project under test. This parsing process involves two steps: first, parsing the abstract syntax tree, and second, performing interprocedural program analysis. The parser extracts key information such as data structures, function implementations, function signatures, and calling relationships. Next, CodeGraphGPT queries API combinations for each library API, focusing on those that have invocation relationships or functional similarities, and generates corresponding fuzz drivers. CodeGraphGPT then attempts to compile these generated fuzz drivers and resolves the compilation errors that arise. During the driver generation and repairing, we provide CodeGraphGPT with a dynamically updated knowledge base on library usage. Once the fuzz drivers are successfully compiled, CodeGraphGPT executes them while monitoring the code coverage of each library file. It employs a code coverage-guided iterative process to mutate API combinations that fail to cover new paths. This iterative process continues until either new code paths are discovered or the mutation budget is exhausted. Finally, CodeGraphGPT uses chain-of-thought reasoning to analyze any crashes produced during fuzzing. To verify the validity of these crashes, we reference an LLM-generated knowledge base containing real CWE-related source code examples.
@@ III-C API Combination Generation
The selection of API combinations is a critical step in guiding the generation of fuzz drivers, as cleverly combining APIs within a single driver can reveal more unexplored execution paths within the target library. CodeGraphGPT employs an external retrieval-augmented process, combining code knowledge graphs with LLM to generate optimal API combinations, which are then used to create fuzz drivers that maximize code coverage. As shown in Figure 1, the CodeGraphGPT framework follows a multi-step method for API combination generation.
@@ III-E Dynamic Program Repair
Due to hallucinations and a lack of high-quality training data, code generated by LLMs often contains syntax and semantic errors [22, 32, 34], which may affect the reliability of fuzz drivers. A robust fuzz driver must ensure correctness at the code level; however, we observed that some fuzz drivers produced by CodeGraphGPT fail to compile. These failures are primarily caused by syntax errors and incorrect usage of library APIs. To address this issue, we developed an LLM-based dynamic program repair mechanism that automatically corrects these errors and improves the overall stability of fuzz drivers, as shown in Algorithm 2. Compared to traditional static rule analysis, LLM-based solutions excel at handling complex, context-dependent errors that are difficult to resolve through static rules alone [33, 15, 16].
@@ 19            $i\leftarrow i+1$
To implement dynamic program repair, we first initialize a knowledge base containing correct library API usages, constructed from samples of fuzz drivers from OSS-Fuzz and the library’s header files. It should be noted that while the OSS-Fuzz code focuses on individual APIs, our approach handles API combinations. The header files help resolve errors caused by missing include statements. When CodeGraphGPT encounters a compilation failure, it processes the compiler error messages and constructs a query to search the external knowledge base for the correct usage of the erroneous API or code snippet. CodeGraphGPT continues to query the knowledge base and apply fixes iteratively until the fuzz driver successfully compiles or a maximum iteration limit is reached.
@@ III-F Fuzzing Loop
The fuzzing loop in CodeGraphGPT is designed to iteratively improve the effectiveness of fuzz testing by initializing high-quality inputs and guiding mutations based on coverage feedback. The loop consists of two main components: the input bank and the coverage-guided mutation mechanism, both working together to maximize code coverage and identify new execution paths.
@@ III-F2 Coverage-guided mutation
During the fuzzing process, CodeGraphGPT monitors the code coverage achieved by each fuzz driver. It analyzes the coverage data across the library and identifies files with lower coverage compared to the overall library. For each low-coverage file, the defined API functions are extracted, and a prioritized list, known as the low coverage API list, is generated based on the coverage rankings of the files.
@@ 23       $i\leftarrow i+1$
This low coverage API list is then used to guide the LLM in mutating and restructuring the current API combinations in the fuzz driver. The objective is to explore new execution paths and increase code coverage by focusing on under-explored APIs. As shown in Algorithm 3, the mutation process is iterative: CodeGraphGPT continues to query the LLM with new API combinations based on the low coverage API list until new branches are covered or computational resources are exhausted.
These mutation strategies are guided by the coverage feedback from previous executions. By prioritizing low-coverage APIs, CodeGraphGPT focuses on areas of the code that are less explored, maximizing the likelihood of discovering new program behaviors.
@@ III-G Crash Analysis
During the fuzzing process, CodeGraphGPT is equipped with several runtime error detection tools, like Address-Sanitizer [35] (ASan), Undefined-Behavior-Sanitizer [25] (UBSan) and MemorySanitizer (MSan) [38]. However, the crashes detected by these tools can originate from different sources, including incorrect API usage in the fuzz driver or bugs within the library APIs themselves. To distinguish between these causes, we developed an automatic crash analysis module designed to identify the root cause of each crash.
@@ IV Evaluation
TABLE I: Code coverage result of fuzz drivers generated by CodeGraphGPT
@@
CodeGraphGPT PromptFuzz OSS-Fuzz
@@
To evaluate the performance of CodeGraphGPT, we defined three research questions:
RQ1: Is CodeGraphGPT more effective than other fuzzers?
RQ2: How does each agent contribute to enhancing the effectiveness of CodeGraphGPT?
RQ3: Can CodeGraphGPT accurately analyze the cause of crashes generated during fuzzing?
@@ IV-A2 Settings
To evaluate CodeGraphGPT, we selected eight open-source libraries as our test subjects. We conducted the experiments using the OSS-Fuzz platform [36] and libFuzzer as the underlying fuzzing environment. We collected the API lists for each library from their respective documentation or official websites. We configured the temperature parameter for DeepSeek-V2-Coder to 0.7 and for DeepSeek-V2-Chat to 1.0. The maximum length for the default API combinations was set to 6, while the maximum number of iterations for program repair was limited to 5. Additionally, the maximum number of mutations for API combinations was set to 3. To ensure that each library was fuzzed for an adequate period, we assigned different time budgets to the fuzz drivers of each library, making sure that the total fuzzing time for each library reached 24 CPU hours. Every experiment was repeated five times to mitigate statistical errors, and the average results were reported.
@@
To address RQ1, we selected eight open-source libraries for evaluation: c-ares, cjson, curl, lcms, libpcap, libtiff, libvpx, and zlib. These libraries were chosen for their diverse API usage and wide adoption in various software projects. We compared CodeGraphGPT with other open-source fuzzers, focusing on coverage-guided fuzzer (OSS-Fuzz [36]) and LLM-based fuzzer (PromptFuzz [27]). During the evaluation, we ran the fuzz drivers from OSS-Fuzz on each library for 24 hours. If a library in OSS-Fuzz contained multiple fuzz drivers, we allocated an equal time budget to each driver to ensure the total fuzzing time was 24 hours. For PromptFuzz, we used the same APIs list provided to CodeGraphGPT and controlled the fuzzing duration to be 24 hours as well. We then evaluated the effectiveness of CodeGraphGPT based on code coverage.
The code coverage results for CodeGraphGPT, PromptFuzz, and OSS-Fuzz are shown in Table I. In comparing CodeGraphGPT with the other fuzzers, CodeGraphGPT demonstrated the highest branch coverage in 6 out of the 8 libraries tested. This superiority can be largely attributed to the structured API combinations generated by CodeGraphGPT, which are derived from a comprehensive code knowledge graph. This allows CodeGraphGPT to capture the complex relationships and semantic dependencies between APIs, thus creating more meaningful fuzz drivers.
In contrast, PromptFuzz initializes its API combinations through random selections and lacks the deeper understanding of API interdependencies. As a result, CodeGraphGPT consistently outperformed PromptFuzz in libraries that require precise API usage patterns. In the lcms library, CodeGraphGPT covered 46.93% of branches, while PromptFuzz covered only 40.71%. These results suggest that CodeGraphGPT’s knowledge-driven approach to API combination allows it to explore a larger portion of the codebase, leading to higher coverage. In cases where CodeGraphGPT did not achieve the highest coverage, such as in the libtiff library, the gap between CodeGraphGPT and PromptFuzz was relatively small (38.81% vs. 48.29%). This suggests that even when PromptFuzz benefits from its constrained input mechanisms, CodeGraphGPT remains competitive, leveraging its knowledge graph to generate targeted fuzz drivers.
@@
In this section, we conduct experiments to investigate the impact of various components of CodeGraphGPT on its effectiveness. The components we evaluate include the code knowledge graph for API combinations generation, compilation repair, and API combination mutation. Detailed experimental results are shown in Figure 2 and Table II.

### v2

# CKGFuzzer: LLM-Based Fuzz Driver Generation Enhanced By Code Knowledge Graph

@@ Abstract
In recent years, the programming capabilities of large language models (LLMs) have garnered significant attention. Fuzz testing, a highly effective technique, plays a key role in enhancing software reliability and detecting vulnerabilities. However, traditional fuzz testing tools rely on manually crafted fuzz drivers, which can limit both testing efficiency and effectiveness. To address this challenge, we propose an automated fuzz testing method driven by a code knowledge graph and powered by an LLM-based intelligent agent system, referred to as CKGFuzzer. We approach fuzz driver creation as a code generation task, leveraging the knowledge graph of the code repository to automate the generation process within the fuzzing loop, while continuously refining both the fuzz driver and input seeds. The code knowledge graph is constructed through interprocedural program analysis, where each node in the graph represents a code entity, such as a function or a file. The knowledge graph-enhanced CKGFuzzer not only effectively resolves compilation errors in fuzz drivers and generates input seeds tailored to specific API usage scenarios, but also analyzes fuzz driver crash reports, assisting developers in improving code quality. By querying the knowledge graph of the code repository and learning from API usage scenarios, we can better identify testing targets and understand the specific purpose of each fuzz driver. We evaluated our approach using eight open-source software projects. The experimental results indicate that CKGFuzzer achieved an average improvement of 8.73% in code coverage compared to state-of-the-art techniques. Additionally, CKGFuzzer reduced the manual review workload in crash case analysis by 84.4% and successfully detected 11 real bugs (including nine previously unreported bugs) across the tested libraries. Our research enhances the overall performance of fuzz testing by refining fuzz driver generation strategies and input seed analysis, offering a more effective solution for vulnerability remediation and software quality improvement.
@@ I Introduction
To address these challenges, we approach fuzz driver generation as a code generation task. By constructing a knowledge graph based on code repositories, we use LLMs to automatically generate fuzz drivers and perform data flow analysis to produce input seeds, thereby achieving fully automated fuzz testing. To enable LLMs to better understand the software under test, we propose abstracting the codebase into a code knowledge graph. In this graph, nodes represent code elements (such as functions, classes, and modules), while edges depict their interactions, with each node containing detailed information about the element’s functionality and behavior. We treat fuzz driver generation as a code planning task, beginning with the formulation of a generation plan followed by iterative fuzz driver generation. To this end, we developed an LLM-driven fuzz driver agent system, CKGFuzzer, based on the code knowledge graph. Our approach fully automates the fuzzing process, significantly reducing engineering complexity while improving the usability of fuzz testing. The code knowledge graph provides deeper insights into code behavior, guiding the LLM in generating higher-quality API combinations for fuzz drivers and coverage-guided mutation.
We evaluated our approach using eight open-source libraries and addressed three research questions (RQs). First, we compared CKGFuzzer with other generation techniques (RQ1). Second, we analyzed the influence of each agent in CKGFuzzer (RQ2). Third, we investigated whether CKGFuzzer can effectively analyze the causes of crashes generated during fuzzing (RQ3). Our experiments demonstrate that CKGFuzzer, supported by the code knowledge graph, can generate high-quality fuzz drivers with enhanced code coverage. Through extensive evaluations across these libraries, we found that the agents of CKGFuzzer play a crucial role in fuzz driver generation. Additionally, our analysis of crash reports indicates that CKGFuzzer can provide valuable insights for developers.
@@
    We propose CKGFuzzer, which leverages a code knowledge graph to enhance fuzz driver generation. By analyzing the relationships between code elements and APIs, CKGFuzzer automatically generates high-quality fuzz drivers that achieve better code coverage.
@@
    CKGFuzzer introduces a fully automated fuzzing framework that integrates multi-agent systems and LLMs. It automates key steps such as fuzz driver generation, input seed initiation, dynamic program repair, and coverage-guided mutation, significantly reducing manual effort and enhancing overall efficiency in the fuzzing process.
@@
    Experimental results show that CKGFuzzer achieved an average increase of 8.73% in code coverage compared to the state-of-the-art methods. Furthermore, CKGFuzzer reduced the manual review workload in crash case analysis by 84.4% and successfully identified 11 real bugs (including nine previously unreported bugs) of tested libraries.
Artifact Availability. We make our artifact public available at https://github.com/Xuhxx/CKGFuzzer.
@@ III-A Overview
Figure 1: Overview of CKGFuzzer
In this section, we present the system design of CKGFuzzer, a fuzz driver generation framework that integrates a multi-agent system with a code knowledge graph. CKGFuzzer is designed to enhance the quality and coverage of fuzz testing by efficiently generating fuzz drivers for API combinations. The primary objective of our framework is to harness the rich and relative information within the code knowledge graph to capture the complex invocation relationships among library APIs and, based on this, generate highly effective fuzz drivers. By introducing a coverage-guided mutation strategy, CKGFuzzer continuously optimizes API combinations, enabling the exploration of additional code paths and the identification of hidden vulnerabilities. Figure 1 illustrates the workflow of CKGFuzzer, which comprises three main components and a unique feedback loop mechanism.
Initially, CKGFuzzer parses the target project and its library APIs, extracting and embedding a code knowledge graph from the project under test. This parsing process involves two steps: first, parsing the abstract syntax tree, and second, performing interprocedural program analysis. The parser extracts key information such as data structures, function implementations, function signatures, and calling relationships. Next, CKGFuzzer queries API combinations for each library API, focusing on those that have invocation relationships or functional similarities, and generates corresponding fuzz drivers. CKGFuzzer then attempts to compile these generated fuzz drivers and resolves the compilation errors that arise. During the driver generation and repairing, we provide CKGFuzzer with a dynamically updated knowledge base on library usage. Once the fuzz drivers are successfully compiled, CKGFuzzer executes them while monitoring the code coverage of each library file. It employs a code coverage-guided iterative process to mutate API combinations that fail to cover new paths. This iterative process continues until either new code paths are discovered or the mutation budget is exhausted. Finally, CKGFuzzer uses chain-of-thought reasoning to analyze any crashes produced during fuzzing. To verify the validity of these crashes, we reference an LLM-generated knowledge base containing real CWE-related source code examples.
@@ III-C API Combination Generation
The selection of API combinations is a critical step in guiding the generation of fuzz drivers, as cleverly combining APIs within a single driver can reveal more unexplored execution paths within the target library. CKGFuzzer employs an external retrieval-augmented process, combining code knowledge graphs with LLM to generate optimal API combinations, which are then used to create fuzz drivers that maximize code coverage. As shown in Figure 1, the CKGFuzzer framework follows a multi-step method for API combination generation.
@@ III-E Dynamic Program Repair
Due to hallucinations and a lack of high-quality training data, code generated by LLMs often contains syntax and semantic errors [22, 32, 34], which may affect the reliability of fuzz drivers. A robust fuzz driver must ensure correctness at the code level; however, we observed that some fuzz drivers produced by CKGFuzzer fail to compile. These failures are primarily caused by syntax errors and incorrect usage of library APIs. To address this issue, we developed an LLM-based dynamic program repair mechanism that automatically corrects these errors and improves the overall stability of fuzz drivers, as shown in Algorithm 2. Compared to traditional static rule analysis, LLM-based solutions excel at handling complex, context-dependent errors that are difficult to resolve through static rules alone [33, 15, 16].
@@ 19            $i\leftarrow i+1$
To implement dynamic program repair, we first initialize a knowledge base containing correct library API usages, constructed from samples of fuzz drivers from OSS-Fuzz and the library’s header files. It should be noted that while the OSS-Fuzz code focuses on individual APIs, our approach handles API combinations. The header files help resolve errors caused by missing include statements. When CKGFuzzer encounters a compilation failure, it processes the compiler error messages and constructs a query to search the external knowledge base for the correct usage of the erroneous API or code snippet. CKGFuzzer continues to query the knowledge base and apply fixes iteratively until the fuzz driver successfully compiles or a maximum iteration limit is reached.
@@ III-F Fuzzing Loop
The fuzzing loop in CKGFuzzer is designed to iteratively improve the effectiveness of fuzz testing by initializing high-quality inputs and guiding mutations based on coverage feedback. The loop consists of two main components: the input bank and the coverage-guided mutation mechanism, both working together to maximize code coverage and identify new execution paths.
@@ III-F2 Coverage-guided mutation
During the fuzzing process, CKGFuzzer monitors the code coverage achieved by each fuzz driver. It analyzes the coverage data across the library and identifies files with lower coverage compared to the overall library. For each low-coverage file, the defined API functions are extracted, and a prioritized list, known as the low coverage API list, is generated based on the coverage rankings of the files.
@@ 23       $i\leftarrow i+1$
This low coverage API list is then used to guide the LLM in mutating and restructuring the current API combinations in the fuzz driver. The objective is to explore new execution paths and increase code coverage by focusing on under-explored APIs. As shown in Algorithm 3, the mutation process is iterative: CKGFuzzer continues to query the LLM with new API combinations based on the low coverage API list until new branches are covered or computational resources are exhausted.
These mutation strategies are guided by the coverage feedback from previous executions. By prioritizing low-coverage APIs, CKGFuzzer focuses on areas of the code that are less explored, maximizing the likelihood of discovering new program behaviors.
@@ III-G Crash Analysis
During the fuzzing process, CKGFuzzer is equipped with several runtime error detection tools, like Address-Sanitizer [35] (ASan), Undefined-Behavior-Sanitizer [25] (UBSan) and MemorySanitizer (MSan) [38]. However, the crashes detected by these tools can originate from different sources, including incorrect API usage in the fuzz driver or bugs within the library APIs themselves. To distinguish between these causes, we developed an automatic crash analysis module designed to identify the root cause of each crash.
@@ IV Evaluation
TABLE I: Code coverage result of fuzz drivers generated by CKGFuzzer
@@
CKGFuzzer PromptFuzz OSS-Fuzz
@@
To evaluate the performance of CKGFuzzer, we defined three research questions:
RQ1: Is CKGFuzzer more effective than other fuzzers?
RQ2: How does each agent contribute to enhancing the effectiveness of CKGFuzzer?
RQ3: Can CKGFuzzer accurately analyze the cause of crashes generated during fuzzing?
@@ IV-A2 Settings
To evaluate CKGFuzzer, we selected eight open-source libraries as our test subjects. We conducted the experiments using the OSS-Fuzz platform [36] and libFuzzer as the underlying fuzzing environment. We collected the API lists for each library from their respective documentation or official websites. We configured the temperature parameter for DeepSeek-V2-Coder to 0.7 and for DeepSeek-V2-Chat to 1.0. The maximum length for the default API combinations was set to 6, while the maximum number of iterations for program repair was limited to 5. Additionally, the maximum number of mutations for API combinations was set to 3. To ensure that each library was fuzzed for an adequate period, we assigned different time budgets to the fuzz drivers of each library, making sure that the total fuzzing time for each library reached 24 CPU hours. Every experiment was repeated five times to mitigate statistical errors, and the average results were reported.
@@
To address RQ1, we selected eight open-source libraries for evaluation: c-ares, cjson, curl, lcms, libpcap, libtiff, libvpx, and zlib. These libraries were chosen for their diverse API usage and wide adoption in various software projects. We compared CKGFuzzer with other open-source fuzzers, focusing on coverage-guided fuzzer (OSS-Fuzz [36]) and LLM-based fuzzer (PromptFuzz [27]). During the evaluation, we ran the fuzz drivers from OSS-Fuzz on each library for 24 hours. If a library in OSS-Fuzz contained multiple fuzz drivers, we allocated an equal time budget to each driver to ensure the total fuzzing time was 24 hours. For PromptFuzz, we used the same APIs list provided to CKGFuzzer and controlled the fuzzing duration to be 24 hours as well. We then evaluated the effectiveness of CKGFuzzer based on code coverage.
The code coverage results for CKGFuzzer, PromptFuzz, and OSS-Fuzz are shown in Table I. In comparing CKGFuzzer with the other fuzzers, CKGFuzzer demonstrated the highest branch coverage in 6 out of the 8 libraries tested. This superiority can be largely attributed to the structured API combinations generated by CKGFuzzer, which are derived from a comprehensive code knowledge graph. This allows CKGFuzzer to capture the complex relationships and semantic dependencies between APIs, thus creating more meaningful fuzz drivers.
In contrast, PromptFuzz initializes its API combinations through random selections and lacks the deeper understanding of API interdependencies. As a result, CKGFuzzer consistently outperformed PromptFuzz in libraries that require precise API usage patterns. In the lcms library, CKGFuzzer covered 46.93% of branches, while PromptFuzz covered only 40.71%. These results suggest that CKGFuzzer’s knowledge-driven approach to API combination allows it to explore a larger portion of the codebase, leading to higher coverage. In cases where CKGFuzzer did not achieve the highest coverage, such as in the libtiff library, the gap between CKGFuzzer and PromptFuzz was relatively small (38.81% vs. 48.29%). This suggests that even when PromptFuzz benefits from its constrained input mechanisms, CKGFuzzer remains competitive, leveraging its knowledge graph to generate targeted fuzz drivers.
@@
In this section, we conduct experiments to investigate the impact of various components of CKGFuzzer on its effectiveness. The components we evaluate include the code knowledge graph for API combinations generation, compilation repair, and API combination mutation. Detailed experimental results are shown in Figure 2 and Table II.